OpenStudy (anonymous):

put this : -x + 3y = -6 into a slope - intercept form.

OpenStudy (anonymous):

y = mx + c m = slope c = y intercept

OpenStudy (anonymous):

ik the form, but i don't know how to put it in the form.

OpenStudy (anonymous):

... what is slope ?

OpenStudy (anonymous):

rearrange the equation into "y=..." then you can compare with the form

OpenStudy (anonymous):

y = m ( -x ) + 3y ?

OpenStudy (anonymous):

noooooo!!!!! you have y in the equation already dont you ? -x + 3y = -6 ^

OpenStudy (anonymous):

rearrange that to get y equal something

OpenStudy (anonymous):

do you know how to rearrange?

OpenStudy (anonymous):

nope. not really :S

OpenStudy (anonymous):

okay, i show you if we have 15x+3y=27 we can put x in the other side so we can have y on it own side without x we do this by doing the opposite to each side... for example +15x+3y=+27 yes? so if we do -15x+15x+3y=+27-15x you see what we do on one side we must on the other the -15x and +15x cancel giving 0 so we are left with 3y=+27-15x we can further simplify this by dividing by 3 giving us y=9-5x do you understand this?

OpenStudy (anonymous):

yea... 80% got this :)

OpenStudy (anonymous):

great, what dont you understand? (20%)?

OpenStudy (anonymous):

we do this by doing the opposite to each side... for example +15x+3y=+27 yes? ( that part )

OpenStudy (anonymous):

okay, the plus and minus cancel out, i only put them there for you to understand, what dont you understand about it?

OpenStudy (anonymous):

ohhh ok nvm i got it.

OpenStudy (mathlegend):

Hello, @Rubie I just wanted to make sure the answer you got for your original problem was correct. Can you tell me what you ended up with? :)

OpenStudy (anonymous):

y = ox + 16/3 ?

OpenStudy (mathlegend):

Okay, no. We have... -x + 3y = -6 We need it to look like: y = mx + b first

OpenStudy (mathlegend):

Can't we move our negative 'x' to the other side of the equal sign by adding it?

OpenStudy (mathlegend):

-x + 3y = -6 +x +x ------------ 3y = x - 6

OpenStudy (anonymous):

so 3y = -6 + -x ?

OpenStudy (anonymous):

ohhh

OpenStudy (mathlegend):

When you move something to the other side the sign changes. So positive would turn negative and a negative would turn positive.

OpenStudy (anonymous):

oooo

OpenStudy (mathlegend):

Just remember I put the 'x' first since it must be in y = mX+b

OpenStudy (anonymous):

ahh

OpenStudy (mathlegend):

3y = x - 6 Now I need to solve for the y. So I must divide by 3 on both sides everything gets divided by this 3 \[\frac{ 3y=x-6 }{ 3 }\]

OpenStudy (mathlegend):

\[y=\frac{ 1 }{ 3 }x-2\]

OpenStudy (mathlegend):

Do you understand? @Rubie

OpenStudy (anonymous):

ohhh thanks for showing me step by step. I get a better visual abt it. thank you :)

OpenStudy (anonymous):

can i ask a bit more to this ?

OpenStudy (mathlegend):

ok

OpenStudy (anonymous):

:| can you identify and explain why ? a) How will the line move Left to Right? b) Steeper or Shallower than the Parent Function?

OpenStudy (mathlegend):

So basically how would I graph this....